# Second- and third-order properties of multidimensional Langevin equations

### v11

@@ 7.3. Inhomogeneous diffusion
Then mathematical induction can be performed when time-reversibility for the second moments holds.
@@
Again, the base case $n=1$ of Eq. (253) is a consequence of Eqs. (245) and (251), while Eq. (252)–(253) can be shown by mathematical induction on $n$.
@@
Once again, the case $n=1$ for Eq. (256) is a consequence of Eqs. (245) and (251), while Eq. (255)–(256) can be shown assuming the case $n=1$.
@@ 7.4. Nonlinear drift
which follows from Eqs. (258)–(259). We also need to use that $\widehat{\mathbf{A}}^{(2)}_{(2)}=\mathbf{A}\otimes\mathbf{A}$ which means that the same rules of time-reversibility that apply to $\mathbf{x}$ and $\mathbf{A}$ also apply to $\mathbf{x}^{(2)}$ and $\widehat{\mathbf{A}}^{(2)}_{(2)}$ (respectively).
@@ 8.1. Linear Gaussian processes: Covariance functions and detailed balance
(295) $\dot{\mathbf{v}}$ $=\mathbf{A}_{\mathbf{x}}\mathbf{x}+\mathbf{A}_{\mathbf{v}}\mathbf{v}+\boldsymbol{\xi},$
(296) $\langle\boldsymbol{\xi}(t){\boldsymbol{\xi}(t^{\prime})}^{\mathsf{T}}\rangle$ $=2\mathbf{D}\delta(t-t^{\prime}),$
where $\boldsymbol{\xi}$ is Gaussian random noise and all coefficients are constant. The first property of interest is:
@@
The quantities that characterize this system are then the position-position covariances $\langle\mathbf{x}\mathbf{x}^{\mathsf{T}}\rangle$, constrained by symmetry, the position-velocity covariances $\langle\mathbf{x}\mathbf{v}^{\mathsf{T}}\rangle$, constrained by antisymmetry, the velocity-velocity covariances $\langle\mathbf{v}\mathbf{v}^{\mathsf{T}}\rangle$, constrained by symmetry, and the quantities $\langle\dot{\mathbf{v}}\mathbf{v}^{\mathsf{T}}\rangle$ interpreted in Itô sense. In place of the latter, we may take the angular momenta $L(v^{i},v^{j})$ together with the diffusivity $\mathbf{D}$.
@@
while Eq. (222) gives a trivial equality, along with:
@@ 8.3. Third-order properties
$\langle x^{i}v^{j}v^{k}\rangle$, $\widetilde{L}(x^{i},v^{j},v^{k})$, $\langle v^{i}v^{j}v^{k}\rangle$, $L(v^{i}v^{j},v^{k})$ (satisfying Eq. (92)), $\langle x^{i}\mathrm{d}[v^{j},v^{k}]/\mathrm{d}t\rangle$, and $\langle v^{i}\mathrm{d}[v^{j},v^{k}]/\mathrm{d}t\rangle$. With the exception of the last two, these quantities can be straightforwardly estimated from time-lapse data without corrections due to fluctuations. The estimation of fluctuation covariance $\mathcal{B}$ and measurement error covariance $\boldsymbol{\Lambda}$ is derived in the Supplementary Material in [10]. However, Eq. (S70) in [10] should read $(6/11)(-1,1,1,-3,1,1)^{\mathsf{T}}$ as in their code. We may write these estimators in terms of the measured values $\mathbf{y}:=\mathbf{x}+\boldsymbol{\eta}$, where $\boldsymbol{\eta}$ is measurement noise. Using the second differences:
@@
Unfortunately, the error in [10] is propagated to the subsequent calculations. Considering inhomogeneous diffusion, Eq. (S77) in [10] should vanish as a result of $\widehat{\mathcal{B}}$ being written in terms of second differences. The next-order correction is obtained by multiplying the force by the measurement error, in place of Eq. (S75). The resulting correction analogous to Eq. (S77) is of order $\mathcal{O}(\Lambda/(\Delta t)^{2})$, which can be ignored because the range of validity of the inference procedure is $\boldsymbol{\Lambda}\ll\langle\mathbf{v}\mathbf{v}^{\mathsf{T}}\rangle(\Delta t)^{2}$. Thus, the choice of position and velocity estimators in Eq. (S92) are of little consequence. The same is true for the position estimator in Eq. (S48). We may choose time-symmetric or -antisymmetric combinations to preserve time-symmetry or -antisymmetry. The error in Eq. (S48) should also be $\mathcal{O}(\Delta t,\Lambda/(\Delta t)^{2})$, where the $\mathcal{O}(\Lambda/(\Delta t)^{2})$ arises from second derivatives with respect to velocity of the basis functions.
@@ 9. Detection of non-Markovianity
(390) $\begin{pmatrix}\dot{x}\\ \dot{y}\end{pmatrix}$ $=\begin{pmatrix}-\lambda&1\\ -\kappa&0\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix}+\boldsymbol{\xi},$
(391) $\langle\boldsymbol{\xi}(t)\rangle$ $=\mathbf{0},$
(392) $\langle\boldsymbol{\xi}(t)\boldsymbol{\xi}^{\mathsf{T}}(t^{\prime})\rangle$ $=2\begin{pmatrix}D&0\\ 0&D^{\prime}\end{pmatrix}\delta(t-t^{\prime}),$
where $\boldsymbol{\xi}$ is Gaussian random noise and all coefficients are constant. For stationarity, we require $\lambda>0$, $\kappa>0$. (We discuss the non-stationary case later.) This formulation allows for feedback $x\to y$, which means that the state of the system could potentially influence the unobserved process. If no such feedback is allowed, then the dynamics instead satisfies:
(393) $\begin{pmatrix}\dot{x}\\ \dot{y}\end{pmatrix}=\begin{pmatrix}-\lambda&1\\ 0&-\eta\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix}+\boldsymbol{\xi},$
for constants $\lambda$, $\eta$, which means that the eigenvalues of the dynamical matrix (called $\mathbf{A}$ in previous sections) are constrained to be real. We will hereafter work with the first formulation as the second is a special case.

### v12

@@ 7.3. Inhomogeneous diffusion
Then mathematical induction can be performed when time-reversibility for the second moments holds. This establishes the time-reversibility of $\langle x^{i}(\tau)x^{j}(0)x^{k}(0)\rangle$.
@@
Again, the base case $n=1$ of Eq. (254) is a consequence of Eqs. (246) and (252), while Eq. (253)–(254) can be shown by mathematical induction on $n$. This establishes the time-reversibility of $\langle\Delta y(\tau)\mathbf{x}(\tau)\mathbf{x}(0)^{\mathsf{T}}\rangle$.
@@
Once again, the case $n=1$ for Eq. (257) is a consequence of Eqs. (246) and (252), while Eq. (256)–(257) can be shown assuming the case $n=1$. This establishes the time-reversibility of $\langle\Delta y(\tau)\mathbf{x}^{(2)}(0)\rangle$.
We may also consider the case of two integrated variables. We will not discuss that here, but we expect similar methods to be applicable and similar results to hold.
@@ 7.4. Nonlinear drift
which follows from Eqs. (259)–(260). We also need to use that $\widehat{\mathbf{A}}^{(2)}_{(2)}=\mathbf{A}\otimes\mathbf{A}$ which means that the same rules of time-reversibility that apply to $\mathbf{x}$ and $\mathbf{A}$ also apply to $\mathbf{x}^{(2)}$ and $\widehat{\mathbf{A}}^{(2)}_{(2)}$ (respectively). This establishes the time-reversibility of $\langle x^{i}(\tau)x^{j}(0)x^{k}(0)\rangle$ with correction $a\mathcal{O}((a,b)^{2})$.
@@ 8.1. Linear Gaussian processes: Covariance functions and detailed balance
(296) $\dot{\mathbf{v}}=\mathbf{A}_{\mathbf{x}}\mathbf{x}+\mathbf{A}_{\mathbf{v}}\mathbf{v}+\bm{\xi},\quad\langle\bm{\xi}(t){\bm{\xi}(t^{\prime})}^{\mathsf{T}}\rangle=2\mathbf{D}\delta(t-t^{\prime}),$
where $\bm{\xi}$ is zero-mean Gaussian white noise and all coefficients are constant. The first property of interest is:
@@
The quantities that characterize this system are then the position-position covariances $\langle\mathbf{x}\mathbf{x}^{\mathsf{T}}\rangle$, constrained by symmetry, the position-velocity covariances $\langle\mathbf{x}\mathbf{v}^{\mathsf{T}}\rangle$, constrained by antisymmetry, the velocity-velocity covariances $\langle\mathbf{v}\mathbf{v}^{\mathsf{T}}\rangle$, constrained by symmetry, and the quantities $\langle\dot{\mathbf{v}}\mathbf{v}^{\mathsf{T}}\rangle$ interpreted in Itô sense. In place of $\langle\dot{\mathbf{v}}\mathbf{v}^{\mathsf{T}}\rangle$, we may take the angular momenta $L(v^{i},v^{j})$ together with the diffusivity $\mathbf{D}$.
@@
while Eq. (223) gives:
@@ 8.3. Third-order properties
$\langle x^{i}v^{j}v^{k}\rangle$, $\widetilde{L}(x^{i},v^{j},v^{k})$, $\langle v^{i}v^{j}v^{k}\rangle$, $L(v^{i}v^{j},v^{k})$ (satisfying Eq. (93)), $\langle x^{i}\mathrm{d}[v^{j},v^{k}]/\mathrm{d}t\rangle$, and $\langle v^{i}\mathrm{d}[v^{j},v^{k}]/\mathrm{d}t\rangle$. With the exception of the last two, these quantities can be straightforwardly estimated from time-lapse data without corrections due to fluctuations. The estimation of fluctuation covariance $\mathcal{B}$ and measurement error covariance $\bm{\Lambda}$ is derived in the Supplementary Material in [10]. However, Eq. (S70) in [10] should read $(6/11)(-1,1,1,-3,1,1)^{\mathsf{T}}$ as in their code. We may write these estimators in terms of the measured values $\mathbf{y}:=\mathbf{x}+\bm{\eta}$, where $\bm{\eta}$ is measurement noise. Using the second differences:
@@
Unfortunately, the error in [10] is propagated to the subsequent calculations. Considering inhomogeneous diffusion, Eq. (S77) in [10] should vanish as a result of $\widehat{\mathcal{B}}$ being written in terms of second differences. The next-order correction is obtained by multiplying the force by the measurement error, in place of Eq. (S75). The resulting correction analogous to Eq. (S77) is of order $\mathcal{O}(\Lambda/(\Delta t)^{2})$, which can be ignored because the range of validity of the inference procedure is $\bm{\Lambda}\ll\langle\mathbf{v}\mathbf{v}^{\mathsf{T}}\rangle(\Delta t)^{2}$. Thus, the choice of position and velocity estimators in Eq. (S92) are of little consequence. The same is true for the position estimator in Eq. (S48). We may choose time-symmetric or -antisymmetric combinations to preserve time-symmetry or -antisymmetry. The error in Eq. (S48) should also be $\mathcal{O}(\Delta t,\Lambda/(\Delta t)^{2})$, where the $\mathcal{O}(\Lambda/(\Delta t)^{2})$ arises from second derivatives with respect to velocity of the basis functions.
@@ 9. Detection of non-Markovianity
(390) $\begin{pmatrix}\dot{x}\\ \dot{y}\end{pmatrix}$ $=\begin{pmatrix}-\lambda&1\\ -\kappa&0\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix}+\bm{\xi},$
(391) $\langle\bm{\xi}(t)\bm{\xi}^{\mathsf{T}}(t^{\prime})\rangle$ $=2\begin{pmatrix}D&0\\ 0&D^{\prime}\end{pmatrix}\delta(t-t^{\prime}),$
where $\bm{\xi}$ is zero-mean Gaussian white noise and all coefficients are constant. For stationarity, we require $\lambda>0$, $\kappa>0$. (We discuss the non-stationary case later.) This formulation allows for feedback $x\to y$, which means that the state of the system could potentially influence the unobserved process. If no such feedback is allowed, then the dynamics instead satisfies:
(392) $\begin{pmatrix}\dot{x}\\ \dot{y}\end{pmatrix}=\begin{pmatrix}-\lambda&1\\ 0&-\eta\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix}+\bm{\xi},$
for constants $\lambda$, $\eta$, which means that the eigenvalues of the matrix of coefficients multiplying the state vector (called $\mathbf{A}$ in previous sections) are constrained to be real. We will hereafter work with the first formulation as the second is a special case.